OpenStudy (pucusana):

1. Solve the linear equation: 15=3x-6

OpenStudy (prettygirl_shynice):

wait i got it

OpenStudy (anonymous):

you can also write it this way: 3x-6=15, you'll still get the same answer. Ok first step, you need to get rid of the 6 in there.

OpenStudy (anonymous):

so now it is 3x = 21

OpenStudy (anonymous):

because you did the opposite of what the six was to cancel it out

OpenStudy (prettygirl_shynice):

step to addd 6 to both sides3x−6+6=15+6 3x=21

OpenStudy (pucusana):

thanks

OpenStudy (prettygirl_shynice):

step 3 dived both sdes by 33x 3 = 21 3 x=7

OpenStudy (prettygirl_shynice):

so ur final answer is 7

OpenStudy (anonymous):

sorry my computer is slow but yeah @Prettygirl_shynice you're right

OpenStudy (prettygirl_shynice):

thx